OpenStudy (anonymous):

4. Find the eleventh term in this arithmetic sequence: .5, .7

OpenStudy (saifoo.khan):

tried it?

OpenStudy (anonymous):

\[d=t1-t2=7-5=2\] \[d=2\] \[a=5\] \[T(n)=a+(n-1)d\] \[=5+(11-1)(2)\] \[=5+10(2)\] \[=5+20\] \[=25\] Am i right?

OpenStudy (anonymous):

.5,.7,.9,.11,.13,.15,.17,.19,.21,.23,.25

OpenStudy (anonymous):

\[T(11)=25\]